OpenStudy (anonymous):

Show how you would solve by Substitution: 2x + y = 0 x – 3y = -7

OpenStudy (anonymous):

Subsitution is where you subsitute one of the variables

OpenStudy (anonymous):

So like if it was x + y = 2 2x + 4y = 2 we would isolate on varaible x + y = 2 y = 2-x and subsitute it x + (2 - x) = 2

OpenStudy (phi):

can you do this problem ?

OpenStudy (anonymous):

im not sure where to even start @phi